# Amortized Safe Active Learning for Real-Time Decision-Making: Pretrained Neural Policies from Simulated Nonparametric Functions

### v1

@@ F.2 Baseline Methods
0:  $\mathcal{D}_{0}\subseteq\mathcal{X}\times\mathcal{Y}$, AL policy $\phi$, $T$
1:  for $t=1,...,T$ do
2:     $\bm{x}_{t}=\phi(T-t+1,\mathcal{D}_{t-1})$ if $\phi$ is budget aware else $\bm{x}_{t}=\phi(\mathcal{D}_{t-1})$
3:     Query at $\bm{x}_{t}$ to get $y_{t}$
4:     $\mathcal{D}_{t}\leftarrow\mathcal{D}_{t-1}\cup\{\bm{x}_{t},y_{t}\}$
5:  end for
6:  return $\mathcal{D}_{T}$ suitable to model $f$
Algorithm A.8 Conventional AL
0:  $\mathcal{D}_{0}\subseteq\mathcal{X}\times\mathcal{Y}$, acquisition function $a$, $T$
1:  for $t=1,...,T$ do
2:     Model $\mathcal{M}_{t-1}$ with $\mathcal{D}_{t-1}$
3:     $\bm{x}_{t}=\text{argmax}_{\bm{x}\in\mathcal{X}}a(x|\mathcal{M}_{t-1},\mathcal{D}_{t-1})$
@@
7:  return $\mathcal{D}_{T}$
Algorithm A.8 is the conventional unconstrained AL (Settles, 2010; Kumar & Gupta, 2020; Tharwat & Schenck, 2023).
Then we write down the safe AL (Schreiter et al., 2015; Zimmer et al., 2018; Li et al., 2022).
@@
0:  $\mathcal{D}_{0}\subseteq\mathcal{X}\times\mathcal{Y}\times\mathcal{Z}$, acquisition function $a$, safety threshold 00, confidence tolerance $\gamma$, $T$
1:  for $t=1,...,T$ do
2:     Model $\mathcal{M}_{t-1}$, $\mathcal{M}_{\text{safety},t-1}$ with $\mathcal{D}_{t-1}$
3:     $\bm{x}_{t}=\text{argmax}_{\bm{x}\in\mathcal{X}}a(x|\mathcal{M}_{t-1},\mathcal{D}_{t-1})$ subject to $p(z(\bm{x})\geq 0|\mathcal{M}_{\text{safety},t-1},\mathcal{D}_{t-1})\geq 1-\gamma$
4:     Query at $\bm{x}_{t}$ to get $y_{t},z_{t}$
5:     $\mathcal{D}_{t}\leftarrow\mathcal{D}_{t-1}\cup\{\bm{x}_{t},y_{t},z_{t}\}$
6:  end for
7:  return $\mathcal{D}_{T}$
@@
For the datasets, the (constrained) acquisition optimizations are solved by an exhaustive search, as done in safe learning literature (Sui et al., 2015; Berkenkamp et al., 2020; Li et al., 2022). In other words, we compute the acquisition scores and the safety distributions on the entire pool of unseen data, and we use the values to solve the (constrained) acquisition optimization.
For function problems, since a fine discretization with exhaustive search is computationally possible, we discretize the space densely by randomly sampling $5000$ input points. Then we solve the (constrained) acquisition optimization problems as if these are pools. Please be aware that the discretization is inherited from conventional safe learning methods (Sui et al., 2015; Berkenkamp et al., 2020; Li et al., 2022)) to solve the constrained acquisition optimization problem. In the main paper, our policies propose queries on continuous space, which requires no discretization.
@@ F.3 Deployment Hardware & Complexities
  • amortized (safe) AL (Algorithms 1 and A.7): NN forward passes take $\mathcal{O}\left((N_{\text{init}}+t-1)^{2}\right)$ at each $t=1,...,T$, dominated by the transformer attender layers (Figure A.B.4Vaswani et al., 2017);
@@ F.4 GP Models
    Our AAL, our ASAL, DAD, Random: we deploy our amortized (safe) AL or the DAD, Random baselines to collect data. After the specified $T$ data points are collected, we use the initial and queried data to fit a GP model with Type II maximum likelihood.
@@
    GP AL, Safe GP AL, Safe Random: We deploy conventional AL and safe AL (Algorithms A.8 and A.9), which update GP iteratively with Type II maximum likelihood.
@@
    AGP AL, Safe AGP AL, Safe ARandom: We deploy conventional AL and safe AL (Algorithms A.8 and A.9), which update GP iteratively with an amortized inference developed by Bitzer et al. (2023) (AGP). Bitzer et al. (2023) sampled GP data and trained a transformer model to approximate the Type II maximum likelihood. This AGP is a model with a transformer module. Whenever an observation dataset is given, the transformer module infers the kernel structure (which we fix to an RBF) together with the kernel hyperparameters. Afterward, one can apply Eq. A.9 in order to compute the predictive distribution. Bitzer et al. (2023) provides a trained model attached to their code.
@@ F.5 Benchmark Problems
$f_{a,b,c,r,s,t}\left((x_{1},x_{2})\right)_{normalize}$ $=\frac{f_{a,b,c,r,s,t}\left((x_{1},x_{2})\right)-mean(f_{a,b,c,r,s,t})}{std(f_{a,b,c,r,s,t})}.$
@@
This is a two dimension dataset described in Rogers et al. (2003) [^4] https://bobby.gramacy.com/surrogates/lgbb.tar.gz, lgbb_original.txt. The dataset has multiple outputs and we take the ”lift” to run our experiments (after normalized to zero mean and unit variance). The inputs are $x_{1}$ (mach) and $x_{2}$ (alpha), which are normalized by
@@
This is a constrained problem (Simionescu, 2014) defined over $(x_{1},x_{2})\in[-1.25,1.25]^{2}$. The main function is
@@
$f\left((x_{1},x_{2})\right)_{normalize}$ $=\frac{f\left((x_{1},x_{2})\right)-mean(f)}{std(f)}.$
@@
$q\left((x_{1},x_{2})\right)_{normalize}$ $=\frac{q\left((x_{1},x_{2})\right)}{std(q)}.$
@@
This is a constrained problem (Townsend, 2017)  [^6] https://www.chebfun.org/examples/opt/ConstrainedOptimization.html defined over $(x_{1},x_{2})\in[-2.25,2.25]\times[-2.5,1.75]$. The main function is
@@
$f\left((x_{1},x_{2})\right)_{normalize}$ $=\frac{f\left((x_{1},x_{2})\right)-mean(f)}{std(f)}.$
@@
$q\left((x_{1},x_{2})\right)_{normalize}$ $=\frac{q\left((x_{1},x_{2})\right)}{std(q)}.$
@@
As described above for unconstrained AL, this dataset has multiple outputs and ”lift” is taken as $y$ We additionally take ”pitch” as $z$ (pitching moment coefficient, see Pamadi et al., 2004). We take a threshold $z\geq 0$, corresponding to around 60% of the space.
@@ G.1 Unconstrained AL Objectives, Main vs Appendix
We compare the two objectives $\mathcal{I},\mathcal{I}_{\text{mean}}$ (Eqs. 5 and A.15) under the same numerical setup as described in Tables A.E.2 and A.F.4. These objectives have the same complexity (Table A.E.3).
@@
G.2 Main Safe AL Objective, $\gamma=0.05$ vs $\gamma=0$
@@
G.3 Amortized Safe AL Objectives Main vs Appendix, Safe AL of Minimum Unsafe Criterion Eq. 6
@@
    conventional GP based safe AL (Algorithm A.9) but we add the unconstrained safety-aware acquisition criterion (Eq. 6), named MinUnsafe GP AL: $\bm{x}_{t}=\text{argmax}\{\mathbb{H}[y(\bm{x})|y_{1:t-1},Y_{\text{init}}]-\log\text{max}(\gamma,p(z(\bm{x})<0|z_{1:t-1},Z_{\text{init}}))\}$ ($\gamma=0.05$, this is the same as Eq. 6 if we take expectation over the forecasted $y(\bm{x})$, and this corresponds to objectives $\mathcal{S}_{\mathcal{H}},\mathcal{S}_{\mathcal{H}_{\text{mean}}}$, see the paragraph of Eqs. 6 and 7).
@@
To avoid confusion, we perform only Type II maximum likelihood for GP modeling. We demonstrate the result on $2D$ problems (Figure A.G.10). Firstly, the performance of Safe GP AL (constrained predictive entropy acquisition) does not seem to have obvious empirical differences from MinUnsafe GP AL. This shows an advantage of MinUnsafe GP AL because its acquisition criterion is safety-aware but unconstrained, and an unconstrained optimization is much easier to be solved. Note that, for a fair comparison, our paper solves the acquisition optimization of MinUnsafe GP AL on discretized set.

### v2

@@ F.2 Baseline methods
1:$\mathcal{D}_{0}\subseteq\mathcal{X}\times\mathcal{Y}$, AL policy $\phi$, $T$
2:for $t=1,...,T$ do
3:     $\bm{x}_{t}=\phi(T-t+1,\mathcal{D}_{t-1})$ if $\phi$ is budget aware else $\bm{x}_{t}=\phi(\mathcal{D}_{t-1})$
@@
7:return $\mathcal{D}_{T}$ suitable to model $f$
Algorithm A.8 Conventional AL
1:$\mathcal{D}_{0}\subseteq\mathcal{X}\times\mathcal{Y}$, acquisition function $a$, $T$
2:for $t=1,...,T$ do
3:     Model $\mathcal{M}_{t-1}$ with $\mathcal{D}_{t-1}$
4:     $\bm{x}_{t}=\text{argmax}_{\bm{x}\in\mathcal{X}}a(x|\mathcal{M}_{t-1},\mathcal{D}_{t-1})$
5:     Query at $\bm{x}_{t}$ to get $y_{t}$
6:     $\mathcal{D}_{t}\leftarrow\mathcal{D}_{t-1}\cup\{\bm{x}_{t},y_{t}\}$
7:end for
8:return $\mathcal{D}_{T}$
Algorithm A.8 is the conventional unconstrained AL [43, 24, 49].
Then we write down the safe AL [40, 55, 26].
@@
1:$\mathcal{D}_{0}\subseteq\mathcal{X}\times\mathcal{Y}\times\mathcal{Z}$, acquisition function $a$, safety threshold 00, confidence tolerance $\gamma$, $T$
2:for $t=1,...,T$ do
3:     Model $\mathcal{M}_{t-1}$, $\mathcal{M}_{\text{safety},t-1}$ with $\mathcal{D}_{t-1}$
4:     $\bm{x}_{t}=\text{argmax}_{\bm{x}\in\mathcal{X}}a(x|\mathcal{M}_{t-1},\mathcal{D}_{t-1})$ subject to $p(z(\bm{x})\geq 0|\mathcal{M}_{\text{safety},t-1},\mathcal{D}_{t-1})\geq 1-\gamma$
5:     Query at $\bm{x}_{t}$ to get $y_{t},z_{t}$
6:     $\mathcal{D}_{t}\leftarrow\mathcal{D}_{t-1}\cup\{\bm{x}_{t},y_{t},z_{t}\}$
7:end for
8:return $\mathcal{D}_{T}$
@@
For the datasets, the (constrained) acquisition optimizations are solved by an exhaustive search, as done in safe learning literature [46, 4, 26]. In other words, we compute the acquisition scores and the safety distributions on the entire pool of unseen data, and we use the values to solve the (constrained) acquisition optimization.
For function problems, since a fine discretization with exhaustive search is computationally possible, we discretize the space densely by randomly sampling $5000$ input points. Then we solve the (constrained) acquisition optimization problems as if these are pools. Please be aware that the discretization is inherited from conventional safe learning methods [46, 4, 26]) to solve the constrained acquisition optimization problem. In the main paper, our policies propose queries on continuous space, which requires no discretization.
@@ F.3 Deployment hardware & complexities
  • amortized (safe) AL (Algorithms 1 and A.7): NN forward passes take $\mathcal{O}\left((N_{\text{init}}+t-1)^{2}\right)$ at each $t=1,...,T$, dominated by the transformer attender layers (Figure A.B.4[52]);
@@ F.4 GP models
    Our AAL, our ASAL, DAD, Random: we deploy our amortized (safe) AL or the DAD, Random baselines to collect data. After the specified $T$ data points are collected, we use the initial and queried data to fit a GP model with Type II maximum likelihood (optimization: L-BFGS-B algorithm).
@@
    GP AL, Safe GP AL, Safe Random: We deploy conventional AL and safe AL (Algorithms A.8 and A.9), which update GP iteratively with Type II maximum likelihood (optimization: L-BFGS-B algorithm).
@@
    AGP AL, Safe AGP AL, Safe ARandom: We deploy conventional AL and safe AL (Algorithms A.8 and A.9), which update GP iteratively with an amortized inference developed by Bitzer et al. [6] (AGP). Bitzer et al. [6] sampled GP data and trained a transformer model to approximate the Type II maximum likelihood. This AGP is a model with a transformer module. Whenever an observation dataset is given, the transformer module infers the kernel structure (which we fix to an RBF) together with the kernel hyperparameters. Afterward, one can apply Eq. A.10 in order to compute the predictive distribution. Bitzer et al. [6] provides a trained model attached to their code.
@@ F.5 Benchmark problems
$f_{a,b,c,r,s,t}\left((x_{1},x_{2})\right)_{\text{normalize}}$ $=\frac{f_{a,b,c,r,s,t}\left((x_{1},x_{2})\right)-mean(f_{a,b,c,r,s,t})}{std(f_{a,b,c,r,s,t})}.$
@@
This is a two dimension dataset described in Rogers et al. [38] [^4] https://bobby.gramacy.com/surrogates/lgbb.tar.gz, lgbb_original.txt. The dataset has multiple outputs and we take the ”lift” to run our experiments (after normalized to zero mean and unit variance). The inputs are $x_{1}$ (mach) and $x_{2}$ (alpha), which are normalized by
@@
This is a constrained problem [44] defined over $(x_{1},x_{2})\in[-1.25,1.25]^{2}$. The main function is
@@
$f\left((x_{1},x_{2})\right)_{\text{normalize}}$ $=\frac{f\left((x_{1},x_{2})\right)-mean(f)}{std(f)}.$
@@
$q\left((x_{1},x_{2})\right)_{\text{normalize}}$ $=\frac{q\left((x_{1},x_{2})\right)}{std(q)}.$
@@
This is a constrained problem [51]  [^6] https://www.chebfun.org/examples/opt/ConstrainedOptimization.html defined over $(x_{1},x_{2})\in[-2.25,2.25]\times[-2.5,1.75]$. The main function is
@@
$f\left((x_{1},x_{2})\right)_{\text{normalize}}$ $=\frac{f\left((x_{1},x_{2})\right)-mean(f)}{std(f)}.$
@@
=$=\begin{cases}\arctan(x_{1}/x_{2})&\text{, if }x_{2}>0\\ \arctan(x_{1}/x_{2})+\text{sign}(x_{1})\pi&\text{, if }x_{2}<0,\text{ say }\text{sign}(0)=1\\ \text{sign}(x_{1})\pi/2&\text{, if }x_{1}\neq 0,x_{2}=0\\ 0&\text{, if }x_{1}=0,x_{2}=0\end{cases},$
$q\left((x_{1},x_{2})\right)_{\text{normalize}}=\frac{q\left((x_{1},x_{2})\right)}{std(q)}.$
@@
As described above for unconstrained AL, this dataset has multiple outputs and ”lift” is taken as $y$. We additionally take ”pitch” as $z$ (pitching moment coefficient, see [33]). We take a threshold $z\geq 0$, corresponding to around 60% of the space.
@@ Appendix G Ablation studies
G.1 Unconstrained AL objectives, main vs appendix
1D appendix $-\mathcal{I}_{\text{mean}}(\phi)$
1D main $-\mathcal{I}(\phi)$
2D appendix $-\mathcal{I}_{\text{mean}}(\phi)$
2D main $-\mathcal{I}(\phi)$
5D appendix $-\mathcal{I}_{\text{mean}}(\phi)$
5D main $-\mathcal{I}(\phi)$
@@ G.1 Unconstrained AL objectives, main vs appendix
We compare the two objectives $\mathcal{I},\mathcal{I}_{\text{mean}}$ (Eqs. 5 and A.16) under the same numerical setup as described in Tables A.E.2 and A.F.4. These objectives have the same complexity (Table A.E.3).
@@
G.2 Unconstrained AL of varied NN sizes
Table A.G.6: Model performance after deploying policies of varied sizes. Shown is the RMSE on the test set (better than random in bold). Smaller NNs (smaller ed) seem to result in noticeable performance deterioration.
AAL ed 128 AAL ed 64 AAL ed 32 GP AL Random
1D: Sinus $0.14\pm 0.004$ $0.13\pm 0.005$ $0.12\pm 0.005$ $0.13\pm 0.009$ $0.33\pm 0.055$
$N_{\text{init}}+T=1+20$
1D: Airline $0.41\pm 0.022$ $0.46\pm 0.027$ $0.42\pm 0.023$ $0.43\pm 0.038$ $0.41\pm 0.023$
$N_{\text{init}}+T=1+20$
2D: Branin $0.21\pm 0.020$ $0.19\pm 0.013$ $0.35\pm 0.032$ $0.19\pm 0.011$ $0.28\pm 0.052$
$N_{\text{init}}+T=1+30$
2D: LGBB $0.17\pm 0.013$ $0.22\pm 0.016$ $0.20\pm 0.014$ $0.17\pm 0.010$ $0.21\pm 0.050$
$N_{\text{init}}+T=1+30$
5D: Airfoil $0.62\pm 0.012$ $0.70\pm 0.017$ $0.70\pm 0.031$ $0.62\pm 0.021$ $0.79\pm 0.047$
$N_{\text{init}}+T=20+40$
As described in Section F.1, the embedding dimension plays a crucial role in the NN size. Here, we would like to study whether we can make the NN even smaller.
We set the embedding dimension, denoted by ed, to 32 or 64, in comparison to our main 128, and we train on the unconstrained AL problems. Similar to the main Table 1, we report the test RMSE after AL deployment. The results are shown in Table A.G.6. It seems that ed 32 can be too small on 2D and 5D problems, ed 64 performs worse than ed 128 on 1D Airline, 2D LGBB, 5D Airfoil but comparable on 1D Sinus and 2D Branin problems. This ablation study shows that our current setting, ed 128, seems to give an effective yet small NNs.
G.3 Main safe AL objective, $\gamma=0.05$ vs $\gamma=0$
@@ G.2 Unconstrained AL of varied NN sizes
G.4 Amortized safe AL objectives main vs appendix, safe AL of minimum unsafe criterion Eq. 7
appendix $\mathcal{S}_{\mathcal{H},\text{division}}(\phi),\mathcal{S}_{\mathcal{H}_{\text{mean}},\text{division}}(\phi)$
main $\mathcal{S}_{\mathcal{H}}(\phi)$, appendix $\mathcal{S}_{\mathcal{H}_{\text{mean}}}(\phi)$, MinUnsafe GP AL
@@
    conventional GP based safe AL (Algorithm A.9) but we add the unconstrained safety-aware acquisition criterion (Eq. 7), named MinUnsafe GP AL: $\bm{x}_{t}=\text{argmax}\{\mathbb{H}[y(\bm{x})|y_{1:t-1},Y_{\text{init}}]-\log\text{max}(\gamma,p(z(\bm{x})<0|z_{1:t-1},Z_{\text{init}}))\}$ ($\gamma=0.05$, this is the same as Eq. 7 if we take expectation over the forecasted $y(\bm{x})$, and this corresponds to objectives $\mathcal{S}_{\mathcal{H}},\mathcal{S}_{\mathcal{H}_{\text{mean}}}$, see the paragraph of Eqs. 7 and 8).
@@
To avoid confusion, we perform only Type II maximum likelihood for GP modeling. We demonstrate the result on all benchmark problems and datasets (Figure A.G.10). Firstly, the performance of Safe GP AL (constrained predictive entropy acquisition) does not seem to have obvious empirical differences from MinUnsafe GP AL. This shows an advantage of MinUnsafe GP AL because its acquisition criterion is safety-aware but unconstrained, and an unconstrained optimization is much easier to be solved. Note that, for a fair comparison, our paper solves the acquisition optimization of MinUnsafe GP AL on discretized set.